# Constraints on Long-Ranged Interactions Between Dark Matter and the Standard Model

### v1

@@ Appendix A Analytic Expressions for Halo Properties
$\sigma_{*}(R_{*})$ $=\sqrt{\frac{(8\sqrt{2}-10)GM_{\chi,{\rm tot}}R_{*}}{3R_{\chi}^{2}}}$ (35)
$\sigma_{\chi}(R_{*})$ $=\sqrt{\frac{GM_{\chi,{\rm tot}}R_{*}}{R_{\chi}^{2}}\ln{\left(\frac{R_{\chi}}{R_{*}}\right)}}$ (36)
@@
$\sigma_{*}(R_{*})$ $\approx\sqrt{\frac{0.12\,GM_{\chi,{\rm tot}}R_{*}}{R_{\chi}^{2}}};$ (37)
the analogous expression for $\sigma_{\chi}(R_{*})$ is extremely unwieldy. These results can be compared to Eqs. (14a) and (14b) in the main text.
We next present the contributions to the self-gravitational potential energy of halos with NFW and $\gamma=1$ profiles, compared to the $\gamma=0$ cases of Eqs. (15b) and (15c). For $\gamma=1$, these are
$U_{*\chi}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&\frac{3}{2}&2\\[2.0pt] 2&\frac{5}{2}&\frac{5}{2}\end{matrix}\right.\right)$ (38a)
$U_{\chi*}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&1&\frac{3}{2}\\[2.0pt] \frac{3}{2}&2&\frac{5}{2}\end{matrix}\right.\right)$ (38b)
@@
$\begin{split}U_{*\chi}\approx\,&\frac{5\,GM_{*,{\rm tot}}M_{\chi,{\rm tot}}}{9R_{*}}\\ &\times\left(\frac{(R_{\chi}^{2}/R_{*})^{2}+\ln(1+R_{\chi}^{2}/R_{*}^{2})}{(1+R_{\chi}^{2}/R_{*}^{2})^{3/2}}\right)\end{split}$ (39a)
$U_{\chi*}\approx\,$ $\frac{0.20\,GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&\frac{3}{2}&2\\[2.0pt] \frac{3}{2}&2&\frac{5}{2}\end{matrix}\right.\right)$ (39b)
@@
For the density overlap form factor, as defined in Eq. (23), we have, for $\gamma=1$,
$\mathcal{F}$ $=\frac{2}{\pi^{5/2}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}0&\frac{1}{2}&1\\[2.0pt] 1&\frac{3}{2}&\frac{5}{2}\end{matrix}\right.\right).$ (40)
@@
$\mathcal{F}$ $\approx 0.032\,G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}\frac{1}{2}&1&1\\[2.0pt] 1&\frac{3}{2}&\frac{5}{2}\end{matrix}\right.\right).$ (41)
@@ Appendix B Heat Transfer with Finite-Range Forces
In this appendix, we generalize the Coulombic heating rate, Eq. (22), to finite-range forces. As in the main text, we take $m_{\chi}\ll m_{*}$ throughout this appendix, as this is the regime relevant to this work.
@@
$\frac{d\sigma}{d\Omega}$ $=4(\alpha Gm_{*}m_{\chi})^{2}\frac{m_{\chi}^{2}}{(\lambda^{-2}+q^{2})^{2}}$ (42)
@@
$\frac{d\sigma}{d\Omega}$ $\lambda\to\infty]{}\frac{\pi}{2}\frac{(\alpha Gm_{*}m_{\chi})^{2}}{(\frac{1}{2}m_{\chi}v_{{\rm rel}}^{2})^{2}(1-\cos\theta)^{2}},$ (43)
@@
$\frac{d\sigma_{T}}{d\Omega}$ $=4(\alpha Gm_{*}m_{\chi})^{2}\frac{m_{\chi}^{2}(1-\cos\theta)}{(\lambda^{-2}+2m_{\chi}^{2}v_{{\rm rel}}^{2}(1-\cos\theta))^{2}}$ (44)
@@
[…]
@@
$\sigma_{T}\xrightarrow[m_{\chi}v_{\rm rel}\lambda\gg 1]{}$ $~{}2\pi{\left(\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\right)}^{2}\ln{\left(\frac{4m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}{1+2\xi m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}\right)}$ (46a)
$\sigma_{T}\xrightarrow[m_{\chi}v_{\rm rel}\lambda\ll 1]{}$ $~{}4\pi\left(4-\xi^{2}\right)\left(\alpha Gm_{*}m_{\chi}^{2}\lambda^{2}\right)^{2}.$ (46b)
@@
While this is not exact, it should be a reasonable approximation for $b_{\rm max}\lesssim\lambda$; excluding $b_{\rm max}\gg\lambda$ should never be significant since the force is exponentially suppressed at such ranges. (Put another way, setting the minimum angle to $\theta>0$, i.e. $\xi>0$, is really only necessary in order to handle situations where $\lambda>R_{*}$ anyway, since otherwise large impact parameters $b>R_{*}$ are exponentially suppressed even without a cutoff.) We can therefore set $b_{\rm max}=\min(\lambda,R_{*})$ without meaningfully affecting any results, and then safely use Eq. (47) with this more conservative $b_{\rm max}$.
@@
$\sigma_{T}\sim$ $2\pi{\left(\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\right)}^{2}\ln\left(4m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}\right)$ for $\xi\ll\frac{1}{m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}$ (48a)
$\sigma_{T}\sim$ $4\pi{\left(\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\right)}^{2}\ln{\left(\frac{v_{\rm rel}^{2}b_{\rm max}}{\alpha Gm_{*}}\right)}$ for $\frac{1}{m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}\ll\xi\ll 2$ (48b)
$\sigma_{T}\sim$ $2\pi b_{\rm max}^{2}$ for $2-\xi\ll 1$ (48c)
@@
$\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\ll\lambda$ (50)
and thus $\sigma_{T}\lesssim\lambda^{2}$ in all three of the cases in Eq. (48) whenever the Born approximation holds.
@@
We can straightforwardly translate Eq. (46a) into a heating rate, as we did for Coulomb scattering in the main text [41, 42]:
$\frac{dU_{\chi}}{dt}\approx\,$ $2\sqrt{2\pi}\alpha^{2}G^{2}\frac{M_{*,{\rm tot}}M_{\chi,{\rm tot}}}{R_{*}^{3}}\mathcal{F}\frac{m_{*}\sigma_{*}^{2}-m_{\chi}\sigma_{\chi}^{2}}{(\sigma_{*}^{2}+\sigma_{\chi}^{2})^{3/2}}\ln{\left(\frac{4m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}}{1+2\xi m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}}\right)}.$ (51)
with $\mathcal{F}$ the form factor defined by Eq. (23), $\xi$ defined by Eq. (47) with $v_{\rm rel}^{2}=\sigma_{*}^{2}+\sigma_{\chi}^{2}$ and $b_{\rm max}=\min(\lambda,R_{*})$, and all of the velocity dispersions evaluated at $R_{*}$. Note that this reproduces the Coulomb result, Eq. (22), if the argument of the log is set to $\Lambda_{G}^{2}$; see Eq. (19). We therefore define for convenience
$\Lambda_{Y}^{2}$ $=\frac{4m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}}{1+2\xi m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}},$ (52)
@@
Acknowledgements.
The authors acknowledge support by NSF Grant PHY-2310429, Simons Investigator Award No. 824870, DOE HEP QuantISED award #100495, the Gordon and Betty Moore Foundation Grant GBMF7946, and the U.S. Department of Energy (DOE), Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DEAC02-07CH11359. ZB is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1656518, by the Dr. HaiPing and Jianmei Jin Fellowship from the Stanford Graduate Fellowship Program, and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab. Fermilab is operated by the Fermi Research Alliance, LLC under Contract DE-AC02-07CH11359 with the U.S. Department of Energy.

### v2

@@ Appendix A Analytic Expressions for Halo Properties
$\sigma_{*}(R_{*})$ $=\sqrt{\frac{(8\sqrt{2}-10)GM_{\chi,{\rm tot}}R_{*}}{3R_{\chi}^{2}}}$ (A.1)
$\sigma_{\chi}(R_{*})$ $=\sqrt{\frac{GM_{\chi,{\rm tot}}R_{*}}{R_{\chi}^{2}}\ln{\left(\frac{R_{\chi}}{R_{*}}\right)}}$ (A.2)
@@
$\sigma_{*}(R_{*})$ $\approx\sqrt{\frac{0.12\,GM_{\chi,{\rm tot}}R_{*}}{R_{\chi}^{2}}};$ (A.3)
the analogous expression for $\sigma_{\chi}(R_{*})$ is extremely unwieldy. These results can be compared to Eqs. (3.8a) and (3.8b) in the main text.
We next present the contributions to the self-gravitational potential energy of halos with NFW and $\gamma=1$ profiles, compared to the $\gamma=0$ cases of Eqs. (3.9b) and (3.9c). For $\gamma=1$, these are
$U_{*\chi}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&\frac{3}{2}&2\\[2.0pt] 2&\frac{5}{2}&\frac{5}{2}\end{matrix}\right.\right)$ (A.4a)
$U_{\chi*}$ $=\frac{4GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{\pi^{3/2}R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&1&\frac{3}{2}\\[2.0pt] \frac{3}{2}&2&\frac{5}{2}\end{matrix}\right.\right)$ (A.4b)
@@
$\begin{split}U_{*\chi}\approx\,&\frac{5\,GM_{*,{\rm tot}}M_{\chi,{\rm tot}}}{9R_{*}}\cdot\frac{(R_{\chi}^{2}/R_{*})^{2}+\ln(1+R_{\chi}^{2}/R_{*}^{2})}{(1+R_{\chi}^{2}/R_{*}^{2})^{3/2}}\end{split}$ (A.5a)
$U_{\chi*}\approx\,$ $\frac{0.20\,GM_{*,{\rm tot}}M_{\chi,{\rm tot}}R_{\chi}^{2}}{R_{*}^{3}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}1&\frac{3}{2}&2\\[2.0pt] \frac{3}{2}&2&\frac{5}{2}\end{matrix}\right.\right)$ (A.5b)
@@
For the density overlap form factor, as defined in Eq. (3.17), we have, for $\gamma=1$,
$\mathcal{F}$ $=\frac{2}{\pi^{5/2}}G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}0&\frac{1}{2}&1\\[2.0pt] 1&\frac{3}{2}&\frac{5}{2}\end{matrix}\right.\right).$ (A.6)
@@
$\mathcal{F}$ $\approx 0.032\,G^{3,3}_{3,3}\left(\frac{R_{*}^{2}}{R_{\chi}^{2}}\left|\begin{matrix}\frac{1}{2}&1&1\\[2.0pt] 1&\frac{3}{2}&\frac{5}{2}\end{matrix}\right.\right).$ (A.7)
@@ Appendix B Heat Transfer with Finite-Range Forces
In this appendix, we generalize the Coulombic heating rate, Eq. (3.16), to finite-range forces. As in the main text, we take $m_{\chi}\ll m_{*}$ throughout this appendix, as this is the regime relevant to this work.
@@
$\frac{d\sigma}{d\Omega}$ $=4(\alpha Gm_{*}m_{\chi})^{2}\frac{m_{\chi}^{2}}{(\lambda^{-2}+q^{2})^{2}}$ (B.1)
@@
$\frac{d\sigma}{d\Omega}$ $\lambda\to\infty]{}\frac{\pi}{2}\frac{(\alpha Gm_{*}m_{\chi})^{2}}{(\frac{1}{2}m_{\chi}v_{{\rm rel}}^{2})^{2}(1-\cos\theta)^{2}},$ (B.2)
@@
$\frac{d\sigma_{T}}{d\Omega}$ $=4(\alpha Gm_{*}m_{\chi})^{2}\frac{m_{\chi}^{2}(1-\cos\theta)}{(\lambda^{-2}+2m_{\chi}^{2}v_{{\rm rel}}^{2}(1-\cos\theta))^{2}}$ (B.3)
@@
$\begin{split}\sigma_{T}=2\pi&{\left(\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\right)}^{2}\left[\ln{\left(\frac{1+4m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}{1+2\xi m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}\right)}+\frac{1}{1+4m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}-\frac{1}{1+2\xi m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}\right].\end{split}$ (B.4)
@@
$\sigma_{T}\xrightarrow[m_{\chi}v_{\rm rel}\lambda\gg 1]{}$ $~{}2\pi{\left(\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\right)}^{2}\ln{\left(\frac{4m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}{1+2\xi m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}\right)}$ (B.5a)
$\sigma_{T}\xrightarrow[m_{\chi}v_{\rm rel}\lambda\ll 1]{}$ $~{}4\pi\left(4-\xi^{2}\right)\left(\alpha Gm_{*}m_{\chi}^{2}\lambda^{2}\right)^{2}.$ (B.5b)
@@
While this is not exact, it should be a reasonable approximation for $b_{\rm max}\lesssim\lambda$; excluding $b_{\rm max}\gg\lambda$ should never be significant since the force is exponentially suppressed at such ranges. (Put another way, setting the minimum angle to $\theta>0$, i.e. $\xi>0$, is only necessary in order to handle situations where $\lambda>R_{*}$ anyway, since otherwise large impact parameters $b>R_{*}$ are exponentially suppressed even without a cutoff.) We can therefore set $b_{\rm max}=\min(\lambda,R_{*})$ without meaningfully affecting any results, and then safely use Eq. (B.6) with this more conservative $b_{\rm max}$.
@@
$\sigma_{T}\sim$ $2\pi{\left(\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\right)}^{2}\ln\left(4m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}\right)$ for $\xi\ll\frac{1}{m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}$ (B.7a)
$\sigma_{T}\sim$ $4\pi{\left(\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\right)}^{2}\ln{\left(\frac{v_{\rm rel}^{2}b_{\rm max}}{\alpha Gm_{*}}\right)}$ for $\frac{1}{m_{\chi}^{2}v_{\rm rel}^{2}\lambda^{2}}\ll\xi\ll 2$ (B.7b)
$\sigma_{T}\sim$ $2\pi b_{\rm max}^{2}$ for $2-\xi\ll 1$ (B.7c)
@@
$\frac{\alpha Gm_{*}}{v_{{\rm rel}}^{2}}\ll\lambda$ (B.9)
and thus $\sigma_{T}\lesssim\lambda^{2}$ in all three of the cases in Eq. (B.7) whenever the Born approximation holds.
@@
We can straightforwardly translate Eq. (B.5a) into a heating rate, as we did for Coulomb scattering in the main text [45, 46, 47]:
$\frac{dU_{\chi}}{dt}\approx\,$ $2\sqrt{2\pi}\alpha^{2}G^{2}\frac{M_{*,{\rm tot}}M_{\chi,{\rm tot}}}{R_{*}^{3}}\mathcal{F}\frac{m_{*}\sigma_{*}^{2}-m_{\chi}\sigma_{\chi}^{2}}{(\sigma_{*}^{2}+\sigma_{\chi}^{2})^{3/2}}\ln{\left(\frac{4m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}}{1+2\xi m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}}\right)}.$ (B.10)
with $\mathcal{F}$ the form factor defined by Eq. (3.17), $\xi$ defined by Eq. (B.6) with $v_{\rm rel}^{2}=\sigma_{*}^{2}+\sigma_{\chi}^{2}$ and $b_{\rm max}=\min(\lambda,R_{*})$, and all of the velocity dispersions evaluated at $R_{*}$. Note that this reproduces the Coulomb result, Eq. (3.16), if the argument of the log is set to $\Lambda_{G}^{2}$; see Eq. (3.13). We therefore define for convenience
$\Lambda_{Y}^{2}$ $=\frac{4m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}}{1+2\xi m_{\chi}^{2}(\sigma_{*}^{2}+\sigma_{\chi}^{2})\lambda^{2}},$ (B.11)
@@
Acknowledgments
The authors acknowledge support by NSF Grant PHY-2310429, Simons Investigator Award No. 824870, DOE HEP QuantISED award #100495, the Gordon and Betty Moore Foundation Grant GBMF7946, and the U.S. Department of Energy (DOE), Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DEAC02-07CH11359. ZB was supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1656518, by the Dr. HaiPing and Jianmei Jin Fellowship from the Stanford Graduate Fellowship Program, and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab. This manuscript has been authored by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.